The Sisyphean Company has a bond outstanding with a face value of $1,000 that reaches maturity in 9 years. The bond certificate indicates that the stated coupon rate for this bond is

8.8​% and that the coupon payments are to be made semiannually. Assuming the appropriate YTM on the Sisyphean bond is

10.3​%, then the price that this bond trades for will be closest​ to:
Answer #1

Information provided:

Face value= future value= $1,000

Time= 9 years*2= 18 semi-annual periods

Coupon rate= 8.8%/2= 4.4%

Coupon payment= 0.044*1,000= $44

Yield to maturity= 10.3%/2= 5.15%

The price of the bond is calculated by computing the present value.

The below has to be entered in a financial calculator to compute the present value:

FV= 1,000

N= 18

PMT= 44

I/Y= 5.15

Press the CPT key and PV to calculate the present value.

The value obtained is 913.35.

Therefore, the price of the bond is $913.35.

In case of any query, kindly comment on the solution.